# Large Language Models for Generative Information Extraction: A Survey

### v2

@@ 4.1 Data Augmentation
Data augmentation involves generating meaningful and diverse data to effectively enhance the training examples or information, while avoiding the introduction of unrealistic, misleading, and offset patterns. Recent powerful LLMs also demonstrate remarkable performance in data generation tasks [211], which has attracted the attention of many researchers using LLMs to generate synthetic data for IE [193, 236, 202, 220, 109, 186, 142]. It can be roughly divided into four strategies according to their techniques, as shown in Fig. 4.
@@
Data Annotation. This strategy directly generates labeled structural data using LLMs. For instance, Zhang et al. [236] proposed LLMaAA to improve accuracy and data efficiency by employing LLMs as annotators within an active learning loop, thereby optimizing both the annotation and training processes. AugURE [202] employed within-sentence pairs augmentation and cross-sentence pairs extraction to enhance the diversity of positive pairs for unsupervised RE, and introduced margin loss for sentence pairs. Li et al. [109] addresses the challenge of document-level relation extraction from a long context, and proposes an automated annotation method for DocRE that combines a LLM with a natural language inference module to generate relation triples.
@@
Knowledge Retrieval. This strategy effectively retrieves related information from LLMs for IE, which is similar to retrieval augmentation generation (RAG) [56]. PGIM [107] presented a two-stage framework for multimodal NER, which leverages ChatGPT as an implicit knowledge base to heuristically retrieve auxiliary knowledge for more efficient entity prediction. Amalvy et al. [4] proposed to improve NER on long documents by generating a synthetic context retrieval training dataset, and training a neural context retriever. Chen and Feng [21] focuses on the task of Multimodal NER and RE, and showcases their approach to enhancing commonsense reasoning skills by employing a range of CoT prompts that encompass different aspects, including nouns, sentences, and multimodal inputs. Additionally, they employ data augmentation techniques such as style, entity, and image manipulation to further improve the performance.
@@
Inverse Generation. This strategy encourages LLMs to generate natural text or questions by utilizing the structural data provided as input, which aligns with the training paradigm of LLMs. For example, SynthIE [85] showed that LLMs can create high-quality synthetic data for complex tasks by reversing the task direction, and train new models that outperformed previous benchmarks. Rather than relying on ground-truth targets, which limits generalizability and scalability, STAR [137] generated structures from valid triggers and arguments, then generates passages with LLMs by designing fine-grained instructions, error identification, and iterative revision. In order to address the challenge of maintaining text coherence while preserving entities, EnTDA [73] proposed a method that involves manipulating the entity list of the original text. This manipulation includes adding, deleting, replacing, or swapping entities. And it further introduces a diversity beam search to enhance diversity in the Entity-to-text generation process.
@@
Overall, these strategies have their own advantages and disadvantages. While data annotation can directly meet task requirements, the ability of LLMs for structured generation still needs improvement. Knowledge retrieval can provide additional information about entities and relations, but it suffers from the hallucination problem and introduces noise. Inverse generation is aligned with the QA paradigm of LLMs. However, it requires structural data and there exists a gap between the generated pairs and the domain that needs to be addressed. Synthetic Datasets for Fine-tuning: While there have been notable advancements in zero-shot scenarios, the process of generating data from ChatGPT and fine-tuning smaller models presents a significant computational resource challenge.
@@ 4.2 Prompt Design
Prompt engineering is a technique employed to enhance the capabilities of LLMs without altering their network parameters [174, 141]. It entails utilizing task-specific instructions, known as prompts, to guide the behavior of model [122, 20]. The practice of prompt design has proven successful in various applications [199]. Undoubtedly, effective prompt design also plays a crucial role in improving the performance of LLMs on IE tasks. In this section, we categorize prompt design approaches based on different strategies and provide a detailed explanation of the underlying motivations behind these techniques:
Question Answer (QA). LLMs are typically instruction-tuned using a dialogue-based method, which creates a gap when compared to the structured prediction requirements of the IE task. Consequently, recent efforts have been made to employ a QA prompt approach to enhance LLMs and facilitate the generation of desired results more seamlessly [233, 210, 70, 156, 100]. For example, QA4RE [233] found that LLMs tend to perform poorly on RE because the instruction-tuning datasets used to train them have a low incidence of RE tasks, and thus proposes reformulating RE as multiple-choice QA to take advantage of the higher prevalence of QA tasks in instruction-tuning datasets. Li et al. [100] analyzed the limitations of existing RE prompts and proposes a new approach called summarize-and-ask prompting, which transforms zero-shot RE inputs into effective QA format using LLMs recursively. It also shows promise in extracting overlapping relations and effectively handling the challenge of none-of-the-above relations. ChatIE [210] proposed a two-stage framework to transform the zero-shot IE task into a multi-turn QA problem. The framework initially identified the different types of elements, and then a sequential IE process is executed for each identified element type. Each stage utilized a multi-turn QA process, where prompts are constructed using templates and previously extracted information.
@@
Self-Improvement. While COT technology can partially elicit the reasoning ability of LLM, it is unavoidable that LLM will still generate factual errors. As a result, there have been efforts [218, 69, 238] to employ LLMs for iterative self-verification and self-improvement, aiming to rectify the results. For instance, Xie et al. [218] proposed a training-free self-improving framework, which consists of three main steps. First, LLMs make predictions on unlabeled corpus, generating self-annotated dataset through self-consistency. Second, the authors explores different strategies to select reliable annotations. Finally, during inference, demonstrations from reliable self-annotated dataset are retrieved for in-context learning. ProgGen [69] involved guiding LLMs to engage in self-reflection within specific domains, resulting in the generation of domain-relevant attributes that contribute to the creation of training data enriched with attributes. Additionally, ProgGen employd a proactive strategy by generating entity terms in advance and constructing NER context data around these entities, thereby circumventing the challenges LLMs face when dealing with intricate structures.
@@ 4.3 Zero-shot Learning
The main challenges in zero-shot learning lie in enabling the model to effectively generalize for tasks and domains that it has not been trained on, as well as aligning the pre-trained paradigm of LLMs. Due to the large amount of knowledge embedded within, LLMs show impressive abilities in zero-shot scenarios of unseen tasks [93, 210]. To achieve zero-shot cross-domain generalization of LLMs in IE tasks, several works have been proposed [175, 249, 204]. These works offered a universal framework for modeling various IE tasks and domains, and introduced innovative training prompts, e.g., instruction [204] and guidelines [175], for learning and capturing the inter-task dependencies of known tasks and generalizing them to unseen tasks and domains. In terms of cross-type generalization, BART-Gen [112] proposed a document-level neural model, by formulating EE task as conditional generation, resulting in better performance and excellent portability on unseen event types.
On the other hand, in order to improve the ability of LLMs under zero shot prompts (no need for further fine-tuning), QA4RE [233] and ChatIE [210] proposed to transform IE into a multi-turn question-answering problem for aligning it with QA task, which is a predominant task in instruction-tuning datasets. Li et al. [100] integrated the chain-of-thought approach and proposed the summarize-and-ask prompting to solve the challenge of ensuring the reliability of outputs from black box LLMs [140].
@@ 4.4 Constrained Decoding Generation
To address this challenge, researchers have explored the use of Constrained generation for better decoding [57, 123, 133, 231]. Constrained decoding generation in autoregressive LLMs refers to the process of generating text while adhering to specific constraints or rules [212, 7, 244]. For example, Geng et al. [57] proposed using grammar-constrained decoding (GCD) as a solution to control the generation of LMs, ensuring that the output follows a given structure. The authors introduced input-dependent grammars to enhance flexibility, allowing the grammar to depend on the input and generate different output structures for different inputs. Unlike previous methods, which generate information token by token, Zaratiana et al. [231] introduced a new approach for extracting entities and relations by generating a linearized graph with nodes representing text spans and edges representing relation triplets. The model used a transformer encoder-decoder architecture with a pointing mechanism and a dynamic vocabulary of spans and relation types, to capture the structural characteristics and boundaries while grounding the output in the original text.
@@ 4.5 Few-shot Learning
Few-shot learning has access to only a limited number of labeled examples, leading to challenges like overfitting and difficulty in capturing complex relationships [76]. Fortunately, scaling up the parameters of LLMs gives them amazing generalization capabilities compared to small pre-trained models, allowing them to achieve excellent performance in few-shot settings [110, 5]. Paolini et al. [160] proposed the Translation between Augmented Natural Languages (TANL) framework; Lu et al. [133] proposed a text-to-structure generation framework (called UIE); and Chen et al. [26] proposed collaborative domain-prefix tuning for NER (called cp-NER). These methods have achieved state-of-the-art performance and demonstrated effectiveness in few-shot setting. Despite the success of LLMs, they face challenges in training-free IE because of the difference between sequence labeling and text-generation models [66]. To overcome these limitations, GPT-NER [203] introduced a self-verification strategy, while GPT-RE [196] enhanced task-aware representations and incorporates reasoning logic into enriched demonstrations. These approaches demonstrate how to effectively leverage the capabilities of GPT for in-context learning. CODEIE [111] and CodeKGC [9] showed that converting IE tasks into code generation tasks with code-style prompts and in-context examples leads to superior performance compared to NL-LLMs. This is because code-style prompts provide a more effective representation of structured output, enabling them to effectively handle the complex dependencies in natural language.
@@ 4.6 Supervised Fine-tuning
Using all training data to fine-tune LLMs is the most common and promising method  [jiang2024toner,cai2024improving,duan2024keyee,liu2024beyond,monajatipoor2024llms,li2024recall,li2024empirical,cai2024improving,li2024relation], which allows the model to capture the underlying structural patterns in the data, and generalize well to unseen samples. For example, DEEPSTRUCT [198] introduced structure pre-training on a collection of task-agnostic corpora to enhance the structural understanding of language models. UniNER [249] explored targeted distillation and mission-focused instruction tuning to train student models for broad applications, such as NER. GIELLM [52] fine-tuned LLMs using mixed datasets, which are collected to utilize the mutual reinforcement effect to enhance performance on multiple tasks.
@@
Astronomical Shao [178] NER ZS Pr GPT-3.5, GPT-4, Claude-2, LLaMA2
@@ 5 Applications on Specific Domains
It is non-ignorable that LLMs have tremendous potential for extracting information from some specific domains, such as mulitmodal [21, 107, 17, 108], multilingual [77, 142, 147], medical [136, 186, 110, 142, 74, 10, 2, 96, 66, 12, 146, 144, 75], scientific [42, 27, 14, 50, 32], astronomical [178, 45], historical [58, 187], and legal [153, 154, 95].
For instance, Chen and Feng [21] introduced a conditional prompt distillation method that enhances a model’s reasoning ability by combining text-image pairs with chain-of-thought knowledge from LLMs, significantly improving performance in multimodal NER and multimodal RE. Tang et al. [186] explored the potential of LLMs in the field of clinical text mining and proposed a novel training approach, which leverages synthetic data to enhance performance and address privacy issues. Dunn et al. [42] presented a sequence-to-sequence approach by using GPT-3 for joint NER and RE from complex scientific text, demonstrating its effectiveness in extracting complex scientific knowledge in material chemistry. Shao [178] explored the use of LLMs to extract astronomical knowledge entities from astrophysical journal articles. Conventional approaches encounter difficulties such as manual labor and limited generalizability. To address these issues, the authors proposed a prompting strategy that incorporates five prompt elements and eight combination prompt, aiming to specifically target celestial object identifiers and telescope names as the experimental objects of interest. González-Gallardo et al. [58] examined the performance of ChatGPT in the NER task specifically on historical texts. The research not only compared ChatGPT with other state-of-the-art language model-based systems but also delved into the challenges encountered in this zero-shot setting. The findings shed light on the limitations of entity identification in historical texts, encompassing concerns related to annotation guidelines, entity complexity, code-switching, and the specificity of prompts.

### v3

@@ 3.2 Relation Extraction
YAYI-UIE [155] SFT $\surd$ Baichuan2-13B - 65.0 - 62.71
KnowCoder [160] SFT $\surd$ LLaMA2-7B - 74.2 - 70.3
@@
3.3 Event Extraction
Events can be defined as specific occurrences or incidents that happen in a given context. Recently, many studies [135, 138] aim to understand events and capture their correlations by extracting event triggers and arguments using LLMs, which is essential for various reasoning tasks [199]. For example, Code4Struct [41] leveraged LLMs to translate text into code to tackle structured prediction tasks, using programming language features to introduce external knowledge and constraints through alignment between structure and code. Considering the interrelation between different arguments in the extended context, PGAD [137] employed a text diffusion model to create a variety of context-aware prompt representations, enhancing both sentence-level and document-level event argument extraction by identifying multiple role-specific argument span queries and coordinating them with context.
3.4 Universal Information Extraction
Code-LLMs. On the other hand, code-based methods unify IE tasks by generating code with a universal programming schema [41]. Code4UIE [6] proposed a universal retrieval-augmented code generation framework, which leverages Python classes to define schemas and uses in-context learning to generate codes that extract structural knowledge from texts. Besides, CodeKGC [159] leveraged the structural knowledge inherent in code and employed the schema-aware prompts and rationale-enhanced generation to improve performance. To enable LLMs to adhere to guidelines out-of-the-box, GoLLIE [32] enhanced zero-shot ability on unseen IE tasks by aligning with annotation guidelines.
3.5 Summaries of Tasks
In this section, we explored the three primary tasks within IE and their associated sub-tasks, as well as frameworks that unify these tasks [4]. A key observation is the increased application of generative LLMs to NER [67, 178], which has seen significant advancements and remains a highly active area of research within IE. In contrast, tasks such as relation extraction and event extraction have seen relatively less application, particularly for strict relation extraction [39] and detection-only event extraction [128]. This discrepancy may be attributed to the critical importance of NER, its utility in various downstream tasks, and its relatively simpler structured outputs, which facilitate large-scale supervised fine-tuning [1].
Additionally, a notable trend is the emergence of unified models for IE tasks, leveraging the general text understanding capabilities of modern large models [4, 156, 6]. Several studies have proposed unified generative frameworks that capture the common abilities across IE tasks and learn the dependencies between them. These unified approaches can be broadly categorized into natural language-based and code-based methods, each with distinct advantages and limitations.
Experimental results summarized in Tables 1, 2, 3 and 4 reveal that universal IE models generally perform better on more complex strict relation extraction tasks due to their ability to learn dependencies across multiple tasks. Furthermore, generative methods significantly outperform discriminative ones in event extraction tasks, particularly in argument classification, highlighting the substantial potential of generative LLMs in this domain.
@@ 4.1 Data Augmentation
Data augmentation involves generating meaningful and diverse data to effectively enhance the training examples, while avoiding the introduction of unrealistic, misleading, and offset patterns. Recent powerful LLMs also demonstrate remarkable performance in data generation tasks [201, 202, 203, 204, 205], which has attracted the attention of many researchers using LLMs to generate synthetic data for IE [127, 61, 101, 44, 161, 162, 163]. It can be roughly divided into four strategies according to their techniques, as shown in Fig. 5.
@@
Data Annotation. This strategy directly generates labeled structural data using LLMs. For instance, Zhang et al. [61] proposed LLMaAA to improve accuracy and data efficiency by employing LLMs as annotators within an active learning loop, thereby optimizing both the annotation and training processes. AugURE [101] employed within-sentence pairs augmentation and cross-sentence pairs extraction to enhance the diversity of positive pairs for unsupervised RE, and introduced margin loss for sentence pairs. Li et al. [161] addressed the challenge of document-level relation extraction from a long context, and proposed an automated annotation method for DocRE that combines a LLM with a natural language inference module to generate relation triples.
@@
Knowledge Retrieval. This strategy effectively retrieves related information from LLMs for IE, which is similar to retrieval augmentation generation (RAG) [206]. PGIM [167] presented a two-stage framework for multimodal NER, which leveraged ChatGPT as an implicit knowledge base to heuristically retrieve auxiliary knowledge for more efficient entity prediction. Amalvy et al. [59] proposed to improve NER on long documents by generating a synthetic context retrieval training dataset, and training a neural context retriever. Chen et al. [166] focused on the task of multimodal NER and RE, and showcased their approach to enhancing commonsense reasoning skills by employing a range of CoT prompts that encompass different aspects, including nouns, sentences, and multimodal inputs. Additionally, they employed data augmentation techniques such as style, entity, and image manipulation to further improve the performance.
@@
Inverse Generation. This strategy encourages LLMs to generate natural text or questions by utilizing the structural data provided as input, which aligns with the training paradigm of LLMs. For example, SynthIE [168] showed that LLMs can create high-quality synthetic data for complex tasks by reversing the task direction, and train new models that outperformed previous benchmarks. Rather than relying on ground-truth targets, which limits generalizability and scalability, STAR [100] generated structures from valid triggers and arguments, then generates passages with LLMs by designing fine-grained instructions, error identification, and iterative revision. In order to address the challenge of maintaining text coherence while preserving entities, EnTDA [58] proposed a method that involves manipulating the entity list of the original text. This manipulation includes adding, deleting, replacing, or swapping entities. And it further introduced a diversity beam search to enhance diversity in the Entity-to-text generation process.
@@ 4.2 Prompt Design
Prompt engineering is a technique employed to enhance the capabilities of LLMs without altering their network parameters [49, 208, 209, 210, 211, 212]. It entails utilizing task-specific instructions, known as prompts, to guide the behavior of model [13, 213, 214]. The practice of prompt design has proven successful in various applications [215, 216, 217, 218]. Undoubtedly, effective prompt design also plays a crucial role in improving the performance of LLMs on IE tasks. In this section, we categorize prompt design approaches based on different strategies and provide a detailed explanation of the underlying motivations behind these techniques:
Question Answer (QA). LLMs are instruction-tuned using a dialogue-based method [219, 220], which creates a gap when compared to the structured prediction requirements of the IE task. Consequently, recent efforts have been made to employ a QA prompt approach to enhance LLMs and facilitate the generation of desired results more seamlessly [98, 40, 90, 108, 96]. For example, QA4RE [98] found that LLMs tend to perform poorly on RE because the instruction-tuning datasets used to train them have a low incidence of RE tasks, and thus proposes reformulating RE as multiple-choice QA to take advantage of the higher prevalence of QA tasks in instruction-tuning datasets. Li et al. [96] analyzed the limitations of existing RE prompts and proposed a new approach called summarize-and-ask prompting, which transforms zero-shot RE inputs into effective QA format using LLMs recursively. It also showed promise in extracting overlapping relations and effectively handling the challenge of none-of-the-above relations. ChatIE [40] proposed a two-stage framework to transform the zero-shot IE task into a multi-turn QA problem. The framework initially identified the different types of elements, and then a sequential IE process is executed for each identified element type. Each stage utilized a multi-turn QA process, where prompts are constructed using templates and previously extracted information.
@@
Self-Improvement. While COT technology can partially elicit the reasoning ability of LLM, it is unavoidable that LLM will still generate factual errors. As a result, there have been efforts [63, 73, 144] to employ LLMs for iterative self-verification and self-improvement, aiming to rectify the results. For instance, Xie et al. [63] proposed a training-free self-improving framework, which consists of three main steps. First, LLMs made predictions on unlabeled corpus, generating self-annotated dataset through self-consistency. Second, the authors explored different strategies to select reliable annotations. Finally, during inference, demonstrations from reliable self-annotated dataset were retrieved for in-context learning. ProgGen [73] involved guiding LLMs to engage in self-reflection within specific domains, resulting in the generation of domain-relevant attributes that contribute to the creation of training data enriched with attributes. Additionally, ProgGen employd a proactive strategy by generating entity terms in advance and constructing NER context data around these entities, thereby circumventing the challenges LLMs face when dealing with intricate structures.
@@ 4.3 Zero-shot Learning
The primary challenges in zero-shot learning involve ensuring that the model can effectively generalize to tasks and domains it has not been trained on, while also aligning the pre-training paradigm of LLMs to these novel tasks. Due to the large amount of knowledge embedded within, LLMs show impressive abilities in zero-shot scenarios of unseen tasks [223, 40]. To achieve zero-shot cross-domain generalization of LLMs in IE tasks, several works have been proposed [32, 64, 5]. These works offered a universal framework for modeling various IE tasks and domains, and introduced innovative training prompts, e.g., instruction [5] and guidelines [32], for learning and capturing the inter-task dependencies of known tasks and generalizing them to unseen tasks and domains. For cross-type generalization, BART-Gen [130] introduced a document-level neural model that frames the EE task as conditional generation, which led to improved performance and strong portability to unseen event types.
On the other hand, in order to improve the ability of LLMs under zero shot prompts (no need for fine-tuning), QA4RE [98] and ChatIE [40] proposed to transform IE into a multi-turn question-answering problem for aligning it with QA task, which is a predominant task in instruction-tuning datasets. Li et al. [96] integrated the chain-of-thought approach and proposed the summarize-and-ask prompting to solve the challenge of ensuring the reliability of outputs from black box LLMs [62].
@@ 4.4 Constrained Decoding Generation
To address this challenge, researchers have explored the use of constrained generation for better decoding [174, 175, 4, 123]. Constrained decoding generation in autoregressive LLMs refers to the process of generating text while adhering to specific constraints or rules [225, 226, 227]. For example, Geng et al. [174] proposed using grammar-constrained decoding as a solution to control the generation of LMs, ensuring that the output follows a given structure. The authors introduced input-dependent grammars to enhance flexibility, allowing the grammar to depend on the input and generate different output structures for different inputs. Unlike previous methods, which generate information token by token, Zaratiana et al. [123] introduced a new approach for extracting entities and relations by generating a linearized graph with nodes representing text spans and edges representing relation triplets. They used a transformer encoder-decoder architecture with a pointing mechanism and a dynamic vocabulary of spans and relation types, to capture the structural characteristics and boundaries while grounding the output in the original text.
@@ 4.5 Few-shot Learning
Few-shot learning has access to only a limited number of labeled examples, leading to challenges like overfitting and difficulty in capturing complex relationships [228]. Fortunately, scaling up the parameters of LLMs gives them amazing generalization capabilities compared to small pre-trained models, allowing them to achieve excellent performance in few-shot settings [91, 43]. Paolini et al. [33] proposed the Translation between Augmented Natural Languages framework; Lu et al. [4] proposed a text-to-structure generation framework; and Chen et al. [60] proposed collaborative domain-prefix tuning for NER. These methods have achieved state-of-the-art performance and demonstrated effectiveness in few-shot setting. Despite the success of LLMs, they face challenges in training-free IE because of the difference between sequence labeling and text-generation models [187]. To overcome these limitations, GPT-NER [42] introduced a self-verification strategy, while GPT-RE [99] enhanced task-aware representations and incorporates reasoning logic into enriched demonstrations. These methods effectively showcase how GPT can be leveraged for in-context learning. CODEIE [36] and CodeKGC [159] showed that converting IE tasks into code generation tasks with code-style prompts and in-context examples leads to superior performance compared to NL-LLMs. This is because code-style prompts provide a more effective representation of structured output, enabling them to effectively handle the complex dependencies in natural language.
@@ 4.6 Supervised Fine-tuning
Using all training data to fine-tune LLMs is the most common and promising method [88, 141, 143, 110, 111, 129, 113, 229, 230, 231, 232, 233], which allows the model to capture the underlying structural patterns in the data, and generalize well to unseen samples. For example, DEEPSTRUCT [151] introduced structure pre-training on a collection of task-agnostic corpora to enhance the structural understanding of language models. UniNER [64] explored targeted distillation and mission-focused instruction tuning to train student models for broad applications, such as NER. GIELLM [34] fine-tuned LLMs using mixed datasets, which are collected to utilize the mutual reinforcement effect to enhance performance on multiple tasks.
4.7 Summaries of Techniques
Data augmentation [61, 101] is a widely explored direction due to its potential in enhancing model performance. LLMs possess extensive implicit knowledge and strong text generation capabilities, making them well-suited for data annotation tasks [222]. However, while data augmentation can expand the training dataset and improve model generalization, they may also introduce noise. For example, knowledge retrieval methods can supply additional context about entities and relationships, enriching the extraction process. Nevertheless, the noise can detract from the overall quality of the extracted information [167, 94].
On the other hand, designing effective prompts remains a significant challenge for leveraging LLMs like GPT-4 [9]. Although approaches such as QA dialogue and CoT [104] strategies can enhance LLMs’ IE capabilities, purely prompt-based methods still lag behind supervised fine-tuning with smaller models. Supervised fine-tuning [64, 67, 5], including cross-domain and few-shot learning, generally yields better performance, which suggests that combining large-scale LLMs for data annotation with supervised fine-tuning using additional data can optimize performance and reduce manual annotation costs [68, 95, 164].
In summary, while various techniques for IE using LLMs offer distinct advantages, they also come with challenges. Thoughtfully combining these strategies can yield significant enhancements in IE tasks.
Table 5: The statistics of research in Specific Domain.
@@ 4.7 Summaries of Techniques
Astronomical Shao et al. [173] NER ZS Pr GPT-3.5, GPT-4, Claude-2, LLaMA2
@@ 5 Applications on Specific Domains
It is non-ignorable that LLMs have tremendous potential for extracting information from some specific domains, such as mulitmodal [166, 167, 57, 94], multilingual [133, 163, 83], medical [183, 162, 91, 163, 184, 171, 185, 186, 187, 188, 85, 179, 172, 234, 235], scientific [180, 181, 80, 47, 182], astronomical [173, 164], historical [189, 126], and legal [89, 78, 115]. Additionally, we present statistical data in Table 5.
For instance, Chen et al. [166] introduced a conditional prompt distillation method that enhances a model’s reasoning ability by combining text-image pairs with chain-of-thought knowledge from LLMs, significantly improving performance in multimodal NER and multimodal RE. Tang et al. [162] explored the potential of LLMs in the field of clinical text mining and proposed a novel training approach, which leverages synthetic data to enhance performance and address privacy issues. Dunn et al. [180] presented a sequence-to-sequence approach by using GPT-3 for joint NER and RE from complex scientific text, demonstrating its effectiveness in extracting complex scientific knowledge in material chemistry. Shao et al. [173] explored the use of LLMs to extract astronomical knowledge entities from astrophysical journal articles. Conventional approaches encounter difficulties such as manual labor and limited generalizability. To address these issues, the authors proposed a prompting strategy that incorporates five prompt elements and eight combination prompt, aiming to specifically target celestial object identifiers and telescope names as the experimental objects of interest. González et al. [189] examined the performance of ChatGPT in the NER task specifically on historical texts. The research not only compared ChatGPT with other state-of-the-art language model-based systems but also delved into the challenges encountered in this zero-shot setting. The findings shed light on the limitations of entity identification in historical texts, encompassing concerns related to annotation guidelines, entity complexity, code-switching, and the specificity of prompts.